Project

General

Profile

20160202 » History » Version 6

Boyd, Steven, 03 February 2016 11:03

1 1 Boyd, Steven
h1. Commissioning and Ops Scheduling meeting : 2/2/2016
2
3
To be held in CR10 in the ATLAS Building @ 15:30 pm 
4
5
Phone access :  ReadyTalk number 
6
US 866-740-1260
7
UK 0800-496-0576
8
Swiss 0800-70-53-51
9
Italy 800182592
10
Access Code: 2155402
11
12
h2. Agenda:
13
14
* Review of tasks and activities identified at last meeting
15
** *Action Item* : John will look at the effect of air on the MC truth scattering angle
16
** *Action Item* : Pierrick/Mark - ensure that the pressure and material in the absorber volumes are recorded 
17
*** *Action Item* : DC to work with Pierrick to ensure this
18
** *Action Item* : Steve to talk to Alan Stevens about stopping 4-hourly monitoring but leaving the alarm in place
19
** *Action Item* : Paolo, Mike and Vicky will investigate the Linde fridge PC instability issue
20
** *Action Item* : Paolo/Pierrick will think of a way to monitor the Linde fridge PC
21
** *Action Item* : Colin, Trevor, Steve, Pierrick to have a meeting to review the Daresbury work schedule in light of running time
22
and the currently known state of the SS work. Meeting to happen on morning of Thursday 28th.
23
24 3 Boyd, Steven
* Current schedule
25
Feb 1 - 5 : Daresbury Electrical work ( 10 A test )
26
Feb 2 : DS coldbox cool down
27
Feb 4 : DS magnet cool down
28
29
30 1 Boyd, Steven
* Updates and future schedule
31
** DS Cooldown schedule
32
** Realignment work and schedule
33
** Cooling water discussion
34
** Controls and Monitoring : Network reconfiguration
35
** Tracker patch panel leak work
36
** Electrical work schedule
37
** Data-taking schedule
38 3 Boyd, Steven
*** User cycle begins Feb 16th
39
*** Start-up activities : Feb 16 - 21
40 4 Boyd, Steven
*** Data-taking starts : Feb 23
41 5 Boyd, Steven
42
h2. Minutes
43
44
Present: Steve B, Chris R, Paul H, Andy N, Vicky B, Mike C, Ryan B, Paul K, Paolo F, Josef F, Henry N, Mark T, Melissa U, Victoria B, John N
45
46
*Action Items
47
** *Action Item* : John will look at the effect of air on the MC truth scattering angle
48
*** _John presented plots of MC scattering angle in air and vacuum. They contain some features that we don't understand. John will continue looking into these._
49
** *Action Item* : Pierrick/Mark - ensure that the pressure and material in the absorber volumes are recorded 
50
*** *Action Item* : DC to work with Pierrick to ensure this
51
*** _Action Item stands_
52
** *Action Item* : Steve to talk to Alan Stevens about stopping 4-hourly monitoring but leaving the alarm in place
53
*** _Email sent but no reply. Steve will try again_
54
** *Action Item* : Paolo, Mike and Vicky will investigate the Linde fridge PC instability issue
55
*** _The IOC needed to be restarted. This fixed the problem_
56
** *Action Item* : Paolo/Pierrick will think of a way to monitor the Linde fridge PC
57
*** _Paolo has added a monitoring process to Nagios to ensure that the IOC on this PC is always alive_
58
** *Action Item* : Colin, Trevor, Steve, Pierrick to have a meeting to review the Daresbury work schedule in light of running time
59
and the currently known state of the SS work. Meeting to happen on morning of Thursday 28th.
60
*** _Action Item postponed. However we do need to do this in the near future. Colin and Trevor spoke informally about priorities and Trevor looked through the schedule. The FC requires about 2 weeks of work. Once it's online some work becomes near-impossible. Trevor is focussing on the FC now. DL are also helping out on the specs of the dump resistors. Work being done on the SS is conditional on the results Steve gets from 10A test. Pierrick reported that the 10A tests are finishing up now. There is no real conclusion on this test. Did find that a small resistance in parallel to the magnet caused overshoots and undershoots so not properly damped. If issues can't be solved we may have to change where the resistors go. Next week most of the PS modification hardware is being sent to RAL so most of the work is done before March is done._
61
*** Issues of chain of authority during "mixed working" i.e. work in rack room at the same time as data-taking needs to be addressed.
62
*** We all agreed that the DC is in overall charge, even during data-taking. However, we need think about the key dance and what needs to
63
be done to ensure clarity in the line of authority.
64
*** *Action Item* : Steve/DC/Andy to think about organisation during mixed work sessions
65
66
67
* Feb 3 DS coldbox cool down
68
* Feb 3 DS magnet cool down
69
70
71
* DS cooldown about to start
72
73
* Realignment work: Drawing for modifications to the bellows are ready to go out. 
74
  Colin doesn't think we can the bellows ready for the data-taking time. Most sensible thing to do is to do as much work as possible but
75
  not move the solenoids now. THis will enable us to take data, before doing the realignment work during the next shutdown period after
76
  user cycle 2015/04. John thinks it's the right way to go. Victoria says that we are not running with fields so the alignment is not
77
  a priority.
78
  Ken : will we have He in the tracker volume.
79
  Colin : we need to know what the leak rate is in both trackers and we should look at them before data-taking.
80
  Next Ops meeting : Leak checking plan. 
81
  Action Item : Colin/Me talk to Melissa about Helium leak fix
82
83
* Cooling water discussion postponed
84
85
* Realignment of PID detectors has been done. Need a survey. Should get the EMR in position. 
86
  * Critical path : what is going on with the tracker. 
87
  * Superconducting permit and PRY discussion : 
88
  Action Item : Andy/Colin/Ken/DC
89
90
* Patch panel work flipped to tracker group
91
92
* Controls and Monitoring. 
93
  All IOCs configured with static IPs. Still need to deploy list of machine/IP number. Each IOC needs to be tested. Network lockout
94
  code is being written. Will be done by the next meeting. 
95
96
* Electrical work in-hand. Electrical test being continued.
97
98
Henry : Test of TOF0, Cerenkov in case magnet flushing had an effect. 
99
100
AOB:
101
Henry: reminds me to see Debbie about phone. Paul K will look at this.
102
Mike: is there any intention to power the FC for the forthcoming User Cycle. When should we put the FC on the frame.
103
Ken : Work plan for the Daresbury team. Action Item : Schedule a meeting between Colin/Trevor/Myself/Pierrick. 
104
Henry : Move to R1. Advance party has gone. Move is not next week, but may be the week after. Colin will set a date for
105
the move. 
106
Pierrick : reminder to Mark that they will look at the SSU Helium.
107
108